OpenStudy (anonymous):

What is the product in simplest form? State any restrictions on the variable.

OpenStudy (anonymous):

\[\frac{ y^2 }{ y-3 }\times \frac{ y^2-y-6 }{ y^2-1y}\]

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

\[\frac{ y(y+2) }{ y+1 }\] This is what I got, but it's not one of the options! Please help for medal and fan :/

OpenStudy (anonymous):

:/ Not sure, I'm sorry for wasting you're time D:

OpenStudy (anonymous):

I'm guessing D... but I need help!